AB INITIO ANHARMONIC FORCE FIELS AND EQUILIBRIUM STRUCTURE OF CARBONYL CHLO-ROFLUORIDE

J.DEMAISON, Laboratoire PhLAM, CNRS, Université de Lille I, Bat. P5, 59655 Villeneuve d’Ascq Cedex, France; \underline{A. PERRIN}, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris Sud, Campus d’Orsay, Bât. 350, 91405 Orsay Cedex, France; H. BÜRGER, FB 9 - Anorganische Chemie - Universität , D-42097 Wuppertal, Germany.

In this work, we report the calculation of a new anharmonic force field up to semi-diagonal quartic terms using a basis set of triple zeta quality. The accuracy of the force field is checked by comparing experimental and ab initio spectroscopic constants. To complete the infrared analysis reported previously ^{a}, the present work involves also the identification of numerous overtone or combination bands of COF^{35}Cl, and for this task new low resolution Fourier transform spectra were recorded at Wuppertal. Finally, two independent equilibrium structures are determined: a purely ab initio one and a semi-experimental one.

aA.Perrin, J.-M.Flaud, H.Burger, G. Pawelke, S.Sander and H. Willner, J. Mol. Spectrosc. \underline{209}, 122 (2001)